OpenStudy (anonymous):

**MEDAL** I'm making flashcards to help me with my ASVAB. Help me with some formulas? Rachel ran ½ mile in 4 minutes. At this rate, how many miles can she run in 15 minutes? A. 1 7/8 <-- B. 4 C. 30 D. 60 Formula?

OpenStudy (nottim):

That's dem military things ain't it?

OpenStudy (anonymous):

Assuming Rachel doesn't get fatigued... If she runs 1/2 mile in 4 minutes, then she is running 1/8 mile every minute (1/8 mile x 4 min = 4/8 = 1/2 mile per 4 min). Now simply multiply the 1/8 mile by the 15 minutes she is running. 1/8 mile x 15 min = 15/8 = 1 and 7/8 or 1.875 miles

OpenStudy (anonymous):

@NotTim The ASVAB? yeah. It's a required test prior to you joining.

OpenStudy (anonymous):

v=d/t where v is the speed in mile/minute, d is the distance in mile, and t is time in minutes. so v=0.5/4=0.125miles/min now you want to find distance. so you rearrange the formula to isolate for d d=vt so now uk v=0.125mile/min and t=15 multiply them together and you get that answer but decimal form.

OpenStudy (anonymous):

Thanks everyone! I wish I could give more then medal.